OpenStudy (anonymous):

Solve radical equation. If there is no solution, write no solution. Uploaded question BELOW (: thanks

OpenStudy (anonymous):

HERE IS THE QUESTION

OpenStudy (anonymous):

2x = sqrt (3x + 1)

OpenStudy (anonymous):

@apoorvk

OpenStudy (anonymous):

I ^2 both sides and then subtracted 3x and 1 from each side. Then I tried factoring but it's unfactorable?? What do I do

OpenStudy (anonymous):

just square both sides

OpenStudy (anonymous):

I did.. then what is the next step?

OpenStudy (anonymous):

\[4x^2 = 3x+1\] then 4x^2 -3x -1=0

OpenStudy (anonymous):

gets??

OpenStudy (anonymous):

then factor

OpenStudy (anonymous):

(4x+1)(x-1)=0

OpenStudy (anonymous):

x = 1

OpenStudy (anonymous):

x=1 and x=-1/4

OpenStudy (anonymous):

@KieraJan did you know how to solve it??

OpenStudy (anonymous):

now I do.. so the answer is one solution?

OpenStudy (anonymous):

because I plug them in to original equation to check, right?

OpenStudy (anonymous):

and when they equal out.. it's considered a solution

OpenStudy (anonymous):

Only one equals out sooo is it one solution???

OpenStudy (anonymous):

I think so